OpenStudy (anonymous):

The population of a country has a relative growth rate of 3% per year. The government is trying to reduce the growth rate to 2%. The population in 1995 was approximately 110 million. Find the projected population for the year 2035 for the following conditions. (Round your answers to the nearest whole number.) (a) The relative growth rate remains at 3% per year. million people (b) The relative growth rate is reduced to 2% per year. million people

OpenStudy (anonymous):

Relative growth rate of a population is solved using the formula: P = N*e^(rt) P = final population N = initial population r = growth rate t = time

OpenStudy (anonymous):

so for a is it P=2035*e^(.03*t)

OpenStudy (anonymous):

i dont get how to put it

OpenStudy (anonymous):

Take the numbers given to you in the problem and figure out what they represent. For example, in the formula, I told you that N = initial population. You plugged in 2035, which is not the initial population. It is the final year. N = initial population.

OpenStudy (anonymous):

so is it 2035=1995*e^(0.03*t)

OpenStudy (anonymous):

No, it's not. The N and the P are not the YEARS for the population, they are the variables for the population itself. I'll give you a hint...you will not plug in the years into the equation at all.

OpenStudy (anonymous):

p=110milion*e^(0.03*T)

OpenStudy (anonymous):

Close. What is t? You can find t using the information given.

OpenStudy (anonymous):

p=110milion*e^(0.03*40)

OpenStudy (anonymous):

Correct. That's A. For B, you need only change the growth rate in the formula.

OpenStudy (anonymous):

p=110milion*e^(0.02*40)

OpenStudy (anonymous):

how am i suppose to put 110 million in the calculator 110,000,000 do I just put that

OpenStudy (anonymous):

Sure, you can. you could also use 1.1 x 10^8

OpenStudy (anonymous):

i put it in the calculator and i get 365212861 for A but it says its wrong

OpenStudy (anonymous):

Try 365,212,862

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

can you help me with B I cant do A I have no more chances

OpenStudy (anonymous):

i put 244809502 for B its wrong

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

See if you can show me your calculations, so I can see.

OpenStudy (anonymous):

i sent a photo

OpenStudy (anonymous):

That should be it. I'm not sure where it's going wrong, but that's the formula and the inputs.

OpenStudy (anonymous):

oh well thanks for the help medal for you !